OpenStudy (anonymous):

QUICK QUESTION 125^b=5 How?

OpenStudy (shubhamsrg):

How? or when ?

OpenStudy (anonymous):

Your question makes my brain hurt

OpenStudy (anonymous):

i think b = 1/3

OpenStudy (anonymous):

How'd you get that answer? (you're right bro)

geerky42 (geerky42):

Use log with base 5

OpenStudy (anonymous):

Eh?

OpenStudy (anonymous):

i dunno, i just knew xD

geerky42 (geerky42):

Remember, \(\log_a b ^x = x\log_a b\), my friend.

geerky42 (geerky42):

Don't you know logarithm?

OpenStudy (anonymous):

no bro, sorry-- not exactly there yet

OpenStudy (anonymous):

i see >.< how about fraction exponents have u been there :D/

OpenStudy (anonymous):

yeth

geerky42 (geerky42):

\(\large a^{\frac{1}{3}} = \sqrt[3]{a}\)

OpenStudy (anonymous):

thats how i got it xD cuz 125^1/3 = cube root of 125.. which is 5 >.<

OpenStudy (anonymous):

OHHHHH RIGHT

OpenStudy (anonymous):

my IMMENSE stupidity amazes me

OpenStudy (anonymous):

Sorry dudes, can't give medals to everyone sooo

OpenStudy (anonymous):

nah, glad i cud help ^.^

OpenStudy (anonymous):

Thanks dudes